(1 point) Consider the probability model with sample space {A,B,C) and P(A-0.1, PB-0.3 PC) 06.Then (a) P(B or C) (b) P(A and B)-
Answer #1

a)P(B or C) =P(B\cup C) = P(B) +P(C)-P(B\cap C)

P(B or C) = 0.3+0.6 -0.3*0.6 = 0.9 -0.18 = 0.72

b)P(A and B) =P(A\cap B) = P(A)*P(B) = 0.1*0.3 = 0.03